OpenStudy (anonymous):

what is the lengths of sides a and b in the 30-60-90 triangle if the hypontenuse is 28 meters

OpenStudy (anonymous):

14 and \[14\sqrt{3}\]

OpenStudy (anonymous):

The side (x) opposite the 30 degree angle satisfies the relationship of sin 30 = x/28. You can get the length of the other side either by the Pythagorean theorem of the cosine relationship.

OpenStudy (anonymous):

*or the cosine relationship. And the sin 30 = 1/2

OpenStudy (anonymous):

thankyou :)

OpenStudy (anonymous):

uw!